OpenStudy (anonymous):

Help please! can someone show me the steps of how to differentiate the following with respect to x 3sin^2x+sec2x

OpenStudy (anonymous):

ddxsin(ax)dx=acos(ax) use this along with chain rule for 3sin^2x

OpenStudy (anonymous):

differentiate each term in turn for 3 sin^2x use the chain rule f'(x) = f'(g(x) * g'(x) also use chain rule for second term

OpenStudy (anonymous):

d/dx [3sin^2(x)] = 6sin x cos x

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Let y=sec 2x dy/dx=d{sec 2x}/dx =sec 2x tan2x×d(2x)/dx USing chain rule =sec 2x tan2x×(2) =2sec 2x tan2x

OpenStudy (anonymous):

@jetly did you use the chain rule to get 6sin x cos x?

OpenStudy (anonymous):

yes kinzan.

OpenStudy (anonymous):

are you happy with the f'(x) notation or do you usually use the Leibnitz fom r kinsan : - dy/dx etc

OpenStudy (anonymous):

*Leibnitz form

OpenStudy (anonymous):

i use the the f'(x) notation but do not get how jetly go 2sec 2x tan2x as his final answer... arnt you meant to use a constant 'u' in the chanirule?

OpenStudy (anonymous):

you can use 'u' for the 'inside function' when using the chain rule - some people find it easier to understand

OpenStudy (anonymous):

i'll differentiate for you using 'u' and leibnitz notation; let y = sec2x and let u = 2x so y = secu dy/du = secu tanu du/dx = 2 so dy/dx = dy/du * du/dx = 2 secu tan u = 2 sec2x tan 2x